# Assessing the Role of Volumetric Brain Information in Multiple Sclerosis Progression

### v1

@@ 2.3 MS progression with classical survival models
To assess the predictive power of the DBM and conventional features, we apply random survival forests (RSF) (Ishwaran et al.,, 2008) to train the tabular dataset on the progression endpoint(s). RSF is an extension of Breiman’s random forest algorithm to the time-to-event setting. In the Appendix, we include additional results using a penalized Cox proportional hazards model with a $\ell_{2}$ penalty (analogous to ridge regression). All analyses in this section were conducted using R (version 4.2.1). RSF models were implemented using the package randomforestSRC (Ishwaran and Kogalur,, 2019) and the Cox model was implemented using the package glmnet (Friedman et al.,, 2021).
@@ 2.4 MS progression with 3D convolutional neural networks
In contrast to the summary statistic featurization of the DBM data, 3D CNNs allow for flexible learning of localized features from the full brain or atlas region voxel-level tensors. Several applied works in MS have considered shallow 3D CNNs (Coll et al.,, 2024; Tousignant et al.,, 2019; Storelli et al.,, 2022). These relatively “parsimonious” architectures can be better suited to tasks with small sample size and without pretrained neural network weights to initialize the model with.
@@
We also train two additional CNN models from the literature, a deep convolutional attention model (CoAtNet), and a hierarchical transformer model designed for medical segmentation tasks (UNesT) (Dai et al.,, 2021; Yu et al.,, 2023). A 3D implementation of CoAtNet is used (Solovyev et al.,, 2022) and it is pretrained using ImageNet (Deng et al.,, 2009). For UNesT, the model is pretrained using T1-w MRIs. For both models, the final layer of their feature extractor modules are fed to an analogous predictor module as the in-house shallow architecture in Figure 3.
@@
3 Results and discussion
@@ 3.1 LOCO-MP feature selection
Region Summary Statistic Full Data Ten Subsamples
Rank $\bar{\Delta}_{j}$ Median Rank Median $\bar{\Delta}_{j}$
@@
In addition to the theoretical guarantees enjoyed by the LOCO-MP framework, we empirically validate its stability by testing its robustness against various data perturbations (Yu,, 2020; Yu and Barter,, 2024). We first assess the robustness of the feature importance list by generating ten 80% subsamples of the data and applying LOCO-MP to each subsample, collecting the corresponding feature importance scores ($\bar{\Delta}_{j}$) and rankings. This is shown in the latter two columns of Table 2. Despite a low signal-to-noise ratio in the DBM data, the top six features are identical between the full data and subsampled data models. The first two features are reversed in rank, but the difference in magnitude is extremely small (less than $0.00003$).
@@
We tested the individual importance of the top DBM features by permuting their values across all patients and evaluating the ranks. For each feature, we permuted its values 25 times across patients and re-applied LOCO-MP for each permutation to evaluate the ranking of the permuted feature. The permutation disrupts the connections between individual DBM features and the outcomes/endpoints and tests whether higher ranks are attributable to specific feature identities. If the ranks of the permuted features are close to those from the unpermuted data, it would imply that the observed importance of the feature is not driven by the uniqueness of the feature (and corresponding brain region) itself. Instead, the perceived feature importance may be due to random and/or other global patterns in the data. The permuted ranks for each feature form a “pseudo-null distribution” under the null hypothesis that the feature is uninformative (shown as a violin plot in Figure 5). This allows us to assess how often permuted ranks are as low or lower than the original rank. The results are shown in Figure 5 – note that “Med” stands for the voxel median feature and “Sd” stands for the voxel standard deviation feature. For the top features identified by LOCO-MP, the ranks from the unpermuted models consistently fall below most, if not all of the permuted ranks, yielding a low “empirical p-value.” This suggests that the feature identities are indeed meaningful and unlikely due to chance, highlighting their potential relevance as medically important regions for MS progression, discussed in the last paragraph of this section.
@@
Figure 5: LOCO permutation test. We plot the distribution of top six LOCO-MP features after permuting their identities and re-applying LOCO-MP to each permutation. The $x$-axis shows the top six features and the $y$-axis shows the rank. “Med” stands for voxel median and “Sd” stands for voxel standard deviation. The green and orange points represent the full data and median subsampled ranks (both unpermuted), respectively. The light blue points represent the rank from a single permutation of that feature.
Benchmarking exercise
@@
To assess the effectiveness of LOCO-MP, we benchmark it against a standard feature selection measurement that ranks features based on their feature importance scores from the random survival forest model. These scores are generated as average difference in out-of-bag prediction error across decision trees when the corresponding feature is permuted. We call this method “RF-Imp.” In principle, the bagging and feature subsets for each decision tree share some similarities with minipatch ensembling. Incorporating shuffled feature values onto already-trained decision trees can create large erroneous changes to predictions that would otherwise have fine performance if retrained without the feature. This may introduce unnecessary instability to RF-Imp when compared with LOCO-MP. Following the same procedure as LOCO-MP, we first compute RF-Imp feature importance scores from the full dataset and then on ten subsamples. Figure 6 shows the top six RF-Imp features based on the full data ranking. The blue boxplot shows the rank distribution across the ten subsamples, and the orange boxplot above it shows the corresponding LOCO-MP subsampled rank distribution for comparison.
@@
Although there is alignment in the set of top regions captured between RF-Imp and LOCO-MP, the RF-Imp results alone demonstrate larger rank instability, especially beyond the third and lateral ventricles. In particular, the 5th and 6th ranked features differ between LOCO-MP and RF-Imp, but the top four are the same. It is clear that LOCO-MP has considerably lower variability in its rank distributions over the ten subsamples, establishing it as a more stable choice for feature selection. In contrast, the RF-Imp ranks have greater spread across subsamples, increasing the possibility of identifying spuriously meaningful features. The greater stability of LOCO-MP allows us to be more confident that the selected features are genuinely meaningful and not due to noise or other artifacts of the data.
@@
Figure 6: Rank distribution for RF-Imp feature importance (blue) vs LOCO-MP (orange). The $y$-axis shows the top six RF-Imp features in terms of their feature importance score from an RSF model applied to the full data. “Med” stands for voxel median and “Sd” stands for voxel standard deviation. The blue boxplots show their subsampled ranks on the $x$-axis. The orange boxplot above to each blue boxplot shows the rank distribution of that feature from LOCO-MP.
@@
In spite of the modestly sized LOCO-MP feature importance scores arising from the small cohort MRI data, the features and their ranks are stable. These features also align with brain regions that are clinically relevant in MS. The identified ventricular regions, including the third and lateral ventricles, are often indicated in MS progression studies (Müller et al.,, 2013; Guenter et al.,, 2022; Simon et al.,, 1999). Moreover, using the same trial data including the treatment arm, Song et al., (2022) identified significant treatment effects in the ventricular regions. The consistency across different studies emphasizes the robustness of the ventricle regions.
The cerebellum is also frequently indicated in MS studies (Weier et al.,, 2015; Brouwer et al.,, 2024), which may play a role in impairment of postural control and balance in MS patients (Gera et al.,, 2020). The precuneus and lingual gyrus are found abnormally activated in the brains of MS patients when performing cognitive exercises, and this activation was later associated with increased mental fatigue and slower speed in task completion (Chen et al.,, 2020). Finally, the parahippocampal gyrus (ranked 6th) has also been linked to varying activation patterns in MS patients compared to controls (Hulst et al.,, 2012), and is also associated with much greater lesion activity compared to controls (Geurts et al.,, 2007).
3.2 Prediction improvements using classical survival models
@@ 3.2.1 CCDP24 outcome
For simplicity, the “Top DBM” models (Top DBM, Conventional + Top DBM) shown in Figure 7 are fit with the top six LOCO-MP features. However, the results from these two feature groupings remained stable with respect to varying numbers of LOCO-MP-selected features. Figure 8 plots the distribution of test-set C-indices from these two feature groupings across the top-ranked DBM features, ranging from 1 to 8. Performance remains stable across different numbers of features, suggesting that the model is robust to how many top features are used – the precise rankings between features and the specific number used for modeling is less important than identifying a broad set of features that can be investigated for clinical relevance. Although these two feature groupings are more investigative, researchers can be confident that the results are not overly sensitive to the specific subset of features. In Section 3.2.2, we discuss extending the LOCO-MP results to an alternative outcome to assess its generalizability.
@@
Despite variations in performance from across different feature sets, the CCDP24 C-indices are modest for even the strongest feature groups. This may arise from various factors, including inherent limitations in the outcome definition which impacts the model’s ability to identify meaningful signal – CCDP24 is characterized by marked changes in multiple events that are sustained for 24 weeks (Montalban et al.,, 2017; Krishnan et al.,, 2023). Although combining multiple events reduces the censoring rate, it comes at the cost of potentially introducing subjective measurements into the analysis: EDSS scoring can vary across studies, and the definition of EDSS-based progression highly depends on the study at hand (Krajnc et al.,, 2021). Therefore, including information from noisy outcome measurements could potentially worsen performance and we hypothesize this may be the case with CCDP24. To combat this, recent literature has pointed to the timed 25-foot walking test as an alternative measure of disability progression that may be more sensitive in detecting genuine disease worsening (Krajnc et al.,, 2021). This test measures the time for a patient to walk 25 feet without any assistive devices. Progression is defined as as a 20% increase in 25-foot walking time that is sustained for 24 weeks (S25FW). Previous studies have demonstrated significantly worse S25FW times in MS patients (Sikes et al.,, 2020) and S25FW has previously been shown to be strongly correlated with other forms of MS progression (Koch et al.,, 2020).
@@
We also note that the two “Top DBM” models in the S25FW analysis use LOCO-MP features identified from the CCDP24 outcome – the algorithm encountered challenges when applied to S25FW due to its high censoring rate, which limited the number of informative minipatches (many minipatches consisted entirely of censored patients). Despite this, the increased performance in the two “Top DBM” S25FW models demonstrates the reliability of LOCO-MP in identifying a stable feature set that captures fundamental aspects of disease progression that are robust across different outcome definitions. This strong generalization is further evidence that the top features may be genuine markers of disease worsening, enhancing its potential for broader clinical applications and underscoring the value of DBM as a regional identification framework.
While this new outcome allowed for more precise patient risk discrimination, the variance of the C-index estimates was higher than that of CCDP24. This is due to the higher censoring rate in S25FW which, by reducing the number of informative events, lowers the effective sample size. This generally leads to increased variance in the estimates because there is less observed data available to accurately capture the underlying risk. Even though S25FW provides clearer risk stratification, the higher variation in the C-index reflects the statistical challenges introduced by a lower effective sample size. In contrast, the CCDP24 outcome has lower variance but lower C-indices as well. It remains an area of future work to create disease measures that are both sensitive and characterized by large sample sizes.
We also observe that the test set C-indices for the strongest feature groupings primarily range between 0.6-0.7, reflecting moderately strong performance at best. These magnitudes are not unexpected given the noisy data regime of the DBM voxels and the low feature importance scores from LOCO-MP, which constrain the predictive power of the DBM features. However, the goal of our study is to evaluate the incremental value of the DBM features as opposed to arguing about the predictive strength of survival models with DBM compared to other MRI imaging features, such as T2 lesion volume or cortical thickness. While DBM features may not outperform these modalities in terms of predictive accuracy, they uniquely capture subtle volumetric abnormalities across the brain and provide orthogonal information to the conventional features. Our models demonstrate how, even in a low-signal regime, DBM features contribute to MS progression in a way that is stable and clinically meaningful, even if their predictive performance may be outperformed by MRI modalities not considered in our study. Future work could explore integrating DBM with well-established MRI modalities to fully assess the strength of each one.
3.3 Prediction improvements using convolutional neural networks
@@
These results underscore the importance of subsetting the DBM input to the top 6 LOCO-MP regions in order to extract signal from a CNN. Each of the full brain CNN models struggle to achieve nonrandom performance (C-index $\geq 0.5$). The shallow full brain CNN features perform modestly better than the complex CoAtNet and UNesT architectures. This suggests that, for this data and cohort size, neither an attention mechanism nor a very deep architecture are able to adequately predict CCDP24. Additionally, CoAtNet was pretrained using ImageNet (Deng et al.,, 2009) and UNesT was pretrained using raw T1-w MRI tensors, indicating that the DBM data may reflect critically different patterns than the raw MRI. Region CNN fares quite poorly when trained on 6 random non-LOCO-MP regions. However, it extracts strong predictive signal when LOCO-MP features are used. The LOCO-MP features were more helpful in the RSF model, achieving a median C-index of 0.61. Region CNN with the top 6 LOCO-MP regions performs comparably on the S25FW outcome (median predictor module C-Index: $0.57$; median RSF retrained C-Index: $0.61$). Ultimately, even the best CNN fares moderately worse than the corresponding classical survival model in Section 3.2, highlighting the challenge of learning complex models for the DBM data on this cohort size. With more data, Region CNN (and CNNs generally) may be able to extract more expressive features from the T1-w MRI data.
4 Conclusion
Identifying predictive and biologically meaningful features in medical studies is an important step towards developing targeted patient interventions and deepening our understanding of disease mechanisms. In this paper, we demonstrate how statistical and machine learning methods can be utilized with DBM to gain meaningful insights about brain regions and their role in predicting MS progression. First, we extend a model-agnostic feature importance method (LOCO-MP) to survival analysis. Using DBM data, we show that LOCO-MP identifies multiple clinically relevant features in a stable manner, outperforming other feature importance metrics. To assess the prognostic value of these features, we fit survival models with various feature groupings that include combinations of conventional and DBM features. We also develop a 3D CNN architecture, Region CNN, that processes a preselected subset of the atlas regions. Our findings indicate that the top DBM features play a valuable role in improving prediction performance, alone or when combined with conventional features. To verify that the selected features can generalize to other outcomes, we tested their performance across alternative outcomes and observed that the models retained strong predictive accuracy.
While our results are specific to this study cohort, the analysis pipeline we develop is general and can be broadly applied to any DBM dataset. Further analyses using external data can develop a more confident scientific understanding about how regional volumetric changes contribute to MS disease worsening. Concerning prediction, our study explored how the DBM features could improve our understanding and prediction of MS progression rather than make definitive claims about model performance. Our findings provide an optimistic and practical benchmark for the prognostic utility of DBM in MS prediction tasks. More complex and principled machine learning frameworks could better contextualize the predictive power of DBM features. Larger cohort sizes of DBM data can additionally enable exploration of whether CNNs can outperform classical approaches, creating the potential for DBM pretrained models.
We evaluated the generalizability of the selected LOCO-MP features by using them to predict a different progression outcome. However, both outcomes were based on 24-week sustained progression measurements which focuses on short-term rather than long-term disease worsening. If the latter is of greater interest, then longitudinal modeling may give insights into whether DBM measurements taken over time are helpful.
Ultimately, our study demonstrates the promise of DBM as a tool for identifying key features and improving MS prediction models. The ability to pinpoint clinically meaningful features in a low-signal regime demonstrates the potential of developing DBM-based biomarkers that can inform therapeutic development and potentially improve efficiency in later MS clinical trials. By integrating DBM with statistical and machine learning frameworks, we provide an analysis pipeline that balances stability, predictive performance, and domain-level interpretability to provide valuable insights to the medical imaging and multiple sclerosis fields.

### v2

@@ 2.3 MS progression with classical survival models
To assess the predictive power of the DBM and conventional features, we apply random survival forests (RSF) [43] to train the tabular dataset on the progression endpoint(s). RSF is an extension of Breiman’s random forest algorithm to the time-to-event setting. In C, we include additional results using a penalized Cox proportional hazards model with a $\ell_{2}$ penalty (analogous to ridge regression) and a $\ell_{1}$ penalty (analogous to lasso regression). All analyses in this section were conducted using R (version 4.2.1). RSF models were implemented using the package randomforestSRC [44] and the Cox models were implemented using the package glmnet [45].
@@ 2.4 MS progression with 3D convolutional neural networks
In contrast to the summary statistic featurization of the DBM data, 3D CNNs allow for flexible learning of localized features from the full brain or atlas region voxel-level tensors. Several applied works in MS have considered shallow 3D CNNs [22, 20, 46]. These relatively “parsimonious” architectures can be better suited to tasks with small sample size and without pretrained neural network weights to initialize the model with.
@@
We also train two additional CNN models from the literature, a deep convolutional attention model (CoAtNet), and a hierarchical transformer model designed for medical segmentation tasks (UNesT) [30, 31]. A 3D implementation of CoAtNet is used [47] and it is pretrained using ImageNet [48]. For UNesT, the model is pretrained using T1-w MRIs. For both models, the final layer of their feature extractor modules are fed to an analogous predictor module as the in-house shallow architecture in Figure 3.
@@
3 Results
We begin by discussing our feature selection results from LOCO-MP and compare this against a traditional random survival forest model (Section 3.1). Sections 3.2 and 3.3 discuss the prediction modeling from traditional survival models and convolutional neural networks, respectively.
@@ 3.1 LOCO-MP feature selection
Region Feature Summary Full Data Ten Subsamples
Statistic Rank $\bar{\Delta}_{j}$ Median Rank Median $\bar{\Delta}_{j}$
@@
In addition to the theoretical guarantees enjoyed by the LOCO-MP framework, we empirically validate its stability by testing its robustness against various data perturbations [50, 51]. We first assess the robustness of the feature importance list by generating ten 80% subsamples of the data and applying LOCO-MP to each subsample, collecting the corresponding feature importance scores ($\bar{\Delta}_{j}$) and rankings. This is shown in the latter two columns of Table 2. Despite a low signal-to-noise ratio in the DBM data, the top six features are identical between the full data and subsampled data models. The first two features are reversed in rank, but the difference in magnitude is extremely small (less than $0.00003$).
@@
We tested the individual importance of the top DBM features by permuting their values across all patients and evaluating the ranks. For each feature, we permuted its values 25 times across patients and re-applied LOCO-MP for each permutation to evaluate the ranking of the permuted feature. The permutation disrupts the connections between individual DBM features and the outcome and tests whether higher ranks are attributable to specific feature identities. If the ranks of the permuted features are close to those from the unpermuted data, it would imply that the observed importance of the feature is not driven by the uniqueness of the feature (and corresponding brain region) itself. Instead, the perceived feature importance may be due to random and/or other global patterns in the data. The permuted ranks for each feature form a “pseudo-null distribution” under the null hypothesis that the feature is uninformative (shown in Figure 5). This allows us to assess how often permuted ranks are as low or lower than the original rank. The results are shown in Figure 5 – note that “Med” stands for the voxel median feature and “Sd” stands for the voxel standard deviation feature. For the top features identified by LOCO-MP, the ranks from the unpermuted models consistently fall below most, if not all of the permuted ranks, yielding a low “empirical p-value.” This suggests that the feature identities are indeed meaningful and unlikely due to chance, highlighting their potential relevance as medically important regions for MS progression, discussed in the last paragraph of this section.
@@
Figure 5: LOCO-MP permutation test. We plot the distribution of the top six LOCO-MP features after 1) permuting their identities and 2) re-applying LOCO-MP to each permutation. The gray histograms show the rank distribution of each of the top LOCO-MP features after it is permuted. “Med” stands for voxel median and “Sd” stands for voxel standard deviation. The blue and red vertical lines represent the full data and median subsampled ranks (both unpermuted), respectively. These lines appear as a single purple line when the permuted and unpermuted ranks are identical.
LOCO-MP vs Random forest-based scoring
To assess the effectiveness of LOCO-MP, we benchmark it against the default feature importance scores produced by the random survival forest model which we refer to as “RF-Imp.” This comparison illustrates the differences between our LOCO-MP approach and the results one might obtain by directly applying standard survival modeling software without first considering the underlying assumptions. While such tools are easily accessible, they may not be tailored for the challenges of high-dimensional feature selection.
@@
The feature importance scores in RF-Imp are generated as the average difference in out-of-bag prediction error across decision trees when the corresponding feature is permuted. In principle, the bagging and feature subsets for each decision tree share some similarities with minipatch ensembling. Incorporating shuffled feature values onto already-trained decision trees can create large erroneous changes to predictions that would otherwise have fine performance if retrained without the feature. This may introduce unnecessary instability to RF-Imp when compared with LOCO-MP. Following the same procedure as LOCO-MP, we first compute RF-Imp feature importance scores from the full dataset and then on ten subsamples. Figure 6 shows the top six RF-Imp features based on the full data ranking. The blue boxplot shows the rank distribution across the ten subsamples, and the orange boxplot above it shows the corresponding LOCO-MP subsampled rank distribution for comparison.
@@
Although there is alignment in the set of top regions captured between RF-Imp and LOCO-MP, the RF-Imp results alone demonstrate larger rank instability, especially beyond the third and lateral ventricles. In particular, the 5th and 6th ranked features differ between LOCO-MP and RF-Imp, but the top four are the same. It is clear that LOCO-MP has considerably lower variability in its rank distributions over the ten subsamples, establishing it as a more stable choice for feature selection. In contrast, the RF-Imp ranks have greater spread across subsamples, increasing the possibility of identifying spuriously meaningful features. As an additional stability check, we replace random forest with gradient boosting machines (GBM) as the underlying survival algorithm in LOCO-MP and find that the subsample ranks of the top regions are concordant with RF, excluding the parahippocampal region (Table 5). The greater stability of LOCO-MP allows us to be more confident that the selected features are genuinely meaningful and not due to noise or other artifacts of the data. In future work, an ensembling of LOCO-MP with multiple survival models may yield an even more stable feature set.
Similar top features are identified when LOCO-MP is combined with survival gradient boosting machines (GBM) instead of random forest, highlighting that LOCO-MP can be relatively algorithm agnostic.
@@
Figure 6: Rank distribution for RF-Imp feature importance (blue) vs LOCO-MP (orange). The $y$-axis shows the top six RF-Imp features in terms of their feature importance score from an RSF model applied to the full data. The $x$-axis plots the ranks across different data subsamples. “Med” stands for voxel median and “Sd” stands for voxel standard deviation. The blue boxplots show their subsampled ranks on the $x$-axis. The orange boxplot above to each blue boxplot shows the rank distribution of that feature from LOCO-MP.
@@
In spite of the modestly sized LOCO-MP feature importance scores arising from the small cohort MRI data, the features and their ranks are stable. These features also align with brain regions that are clinically relevant in MS. The two identified ventricular regions, including the third and lateral ventricles, are often indicated in MS progression studies [52, 53, 54]. The consistency of ventricle-related findings across different studies emphasizes the robustness of these regions. Notably, these regions were also identified as significant in a recent causal study using the treatment arm of the same trial [6]. While this study focused on treatment effects, our use of the control arm to predict progression confirms the continued importance of the ventricular regions even in the absence of therapeutic intervention.
The cerebellum is also frequently indicated in MS studies [55, 56], which may play a role in impairment of postural control and balance in MS patients [57]. The precuneus and lingual gyrus are found abnormally activated in the brains of MS patients when performing cognitive exercises, and this activation was later associated with increased mental fatigue and slower speed in task completion [58]. Finally, the parahippocampal gyrus (ranked 6th) has also been linked to varying activation patterns in MS patients compared to controls [59], and is also associated with much greater lesion activity compared to controls [60].
While the identified regions possess some degree of clinical relevance in previous MS studies, we emphasize that our study is exploratory in this area and does not intend to draw any clinical conclusions. Pathological development of MS can be diffusive across brain anatomy and functions [1, 61]. As a result, many brain regions can atrophy or be implicated by the disease. In order to inform future clinical decision-making, further validation on larger and more heterogeneous patient populations is required. However, our study provides a practical first step towards future scientific studies on certain brain regions.
3.2 Prediction improvements from selected regions under classical survival models
@@ 3.2.1 CCDP24 outcome
For simplicity, the “Top DBM” models (Top DBM, Conventional + Top DBM) shown in Figure 7 are fit with the top six LOCO-MP features. However, the results from these two feature groupings remained stable with respect to varying numbers of LOCO-MP-selected features. Figure 8 plots the distribution of test-set C-indices from these two feature groupings across the top-ranked DBM features, ranging from 1 to 8. Performance remains stable across different numbers of features, suggesting that the model is robust to how many top features are used – the precise rankings between features and the specific number used for modeling is less important than identifying a broad set of features that can be investigated for clinical relevance. Although these two feature groupings are more investigative, researchers can be confident that the results are not overly sensitive to the specific subset of features. To assess generalizability of the Top DBM features, Section 3.2.2 applies this feature set to an alternative progression outcome, and Section D applies this feature set to a separate control cohort.
@@
Despite variations in performance from across different feature sets, the CCDP24 C-indices are modest for even the strongest feature groups. This may arise from various factors, including inherent limitations in the outcome definition which impacts the model’s ability to identify meaningful signal – CCDP24 is characterized by marked changes in multiple events that are sustained for 24 weeks [62, 63]. Although combining multiple events reduces the censoring rate, it comes at the cost of potentially introducing subjective measurements into the analysis: EDSS scoring can vary across studies, and the definition of EDSS-based progression highly depends on the study at hand [64]. Therefore, including information from noisy outcome measurements could potentially worsen performance and we hypothesize this may be the case with CCDP24. To combat this, recent literature has pointed to the timed 25-foot walking test as an alternative measure of disability progression that may be more sensitive in detecting genuine disease worsening [64]. This test measures the time for a patient to walk 25 feet without any assistive devices. Progression is defined as a 20% increase in 25-foot walking time that is sustained for 24 weeks (S25FW). Previous studies have demonstrated significantly worse S25FW times in MS patients [65] and S25FW has previously been shown to be strongly correlated with other forms of MS progression [66].
@@
We also note that the two “Top DBM” models in the S25FW analysis use LOCO-MP features identified from the CCDP24 outcome – the algorithm encountered challenges when applied to S25FW due to its high censoring rate, which limited the number of informative minipatches (many minipatches consisted entirely of censored patients). Despite this, the increased performance in the two “Top DBM” S25FW models demonstrates the reliability of LOCO-MP in identifying a stable feature set that captures fundamental aspects of disease progression that are robust across different outcome definitions. D further demonstrates that these identified DBM improve S25FW prediction in a separate validation cohort. These strong generalizations are further evidence that the top features may be genuine markers of disease worsening, enhancing its potential for broader clinical applications and underscoring the value of DBM as a regional identification framework.
While this new outcome allowed for more precise patient risk discrimination, the variance of the C-index estimates was higher than that of CCDP24. This is due to the higher censoring rate in S25FW which, by reducing the number of informative events, lowers the effective sample size. This generally leads to increased variance in the estimates because there is less observed data available to accurately capture the underlying risk. Even though S25FW provides clearer risk stratification, the higher variation in the C-index reflects the statistical challenges introduced by a lower effective sample size. In contrast, the CCDP24 outcome has lower variance but lower C-indices as well.
3.3 Prediction improvements from selected regions under convolutional neural networks
@@
These results underscore the importance of subsetting the DBM input to the top 6 LOCO-MP regions in order to extract signal from a CNN. Each of the full brain CNN models struggle to achieve nonrandom performance (C-index $\geq 0.5$). The shallow full brain CNN features perform modestly better than the complex CoAtNet and UNesT architectures. This suggests that, for this data and cohort size, neither an attention mechanism nor a very deep architecture are able to adequately predict CCDP24. Additionally, CoAtNet was pretrained using ImageNet [48] and UNesT was pretrained using raw T1-w MRI tensors, indicating that the DBM data may reflect critically different patterns than the raw MRI. Region CNN fares quite poorly when trained on 6 random non-LOCO-MP regions. However, it extracts strong predictive signal when LOCO-MP features are used. The LOCO-MP features were more helpful in the RSF model, achieving a median C-index of 0.61. Region CNN with the top 6 LOCO-MP regions performs comparably on the S25FW outcome (median predictor module C-Index: $0.57$; median RSF retrained C-Index: $0.61$). Ultimately, even the best CNN fares moderately worse than the corresponding classical survival model in Section 3.2, highlighting the challenge of learning complex models for the DBM data on this cohort size. With more data, Region CNN (and CNNs generally) may be able to extract more expressive features from the T1-w MRI data.
4 Discussion
This study develops a data analysis pipeline that balances stability, predictive performance, and domain-level interpretability to improve MS prediction models with DBM-based features. We demonstrate how LOCO-MP identifies clinically relevant features in a stable manner, outperforming other feature importance approaches. Our sample size of 350 is modest for complex prediction tasks such as MS progression prediction from neuroimaging data. While sample size can influence the stability and generalizability of model outputs, we mitigated these concerns by conducting extensive stability checks discussed in Section 3.1 and varying the number of top DBM features used for modeling (Figures 8 and 10). The selected LOCO-MP features generalized well to a separate validation cohort (D). We note that our analysis pipeline is scalable and well-suited to larger datasets as they become available.
Regarding prediction performance, we observe that the test set C-indices for the strongest feature groupings primarily range between 0.65-0.75. We note that our prediction task focuses on 24-week progression, a relatively short time frame for a chronic illness such as MS. Predicting over such brief periods is inherently challenging, as disease progression tends to be subtler and more variable, limiting the ability of models to capture robust patterns. This limitation likely prevents the observed C-index from being higher. In fact, several other RRMS patient studies that predict short-term progression with similar sample sizes report comparable C-indices or AUROC values [^6] The C-index is analogous to the AUROC in binary prediction tasks with survival data. (between 0.48 and 0.73) when using a similarly-defined outcome with different types of brain data [18, 67, 68, 69]. Another study reported slightly higher C-indices ($0.76$) when integrating genotypic factors with clinical variables [70] – the use of non-MRI based features, such as those based on genetic or gut microbiome information, has proven to yield strong predictions and is a potential future direction for MS prediction [71]. In addition, Pellegrini et al. (2020) observed that C-indices did not exceed 0.65 with conventional survival models, despite using a larger cohort [18].
The endpoints themselves also have limitations. As mentioned in Section 3.2.2, CCDP24 is derived by combining the scores of EDSS, 9HPT, and T25FW. While each of the three clinical assessments has its own limitations [72, 73, 74], EDSS is more subjective and less sensitive to disease-related changes [72]. S25FW is a quantitative measure of lower limb movement function. Because it only captures one specific function, its two-year progression rate experiences more censoring. Progression on this metric regardless has a high correspondence with overall RRMS progression [73]. It remains an area of future work to find a clinical progression measure that optimally balances sensitivity and censoring rates.
We additionally explored the performance of 3D CNN models using the full brain versus the LOCO-MP top regions. Though the 3D CNN models substantially improved when constraining the DBM input to the LOCO-MP top DBM regions, their performance still lagged behind conventional modeling using summary statistics of these regions. State of the art 3D CNN architectures did not change this story; the small sample size could be the main limiting factor for deep learning performance. In the future, we are interested in a deeper exploration of small sample size strategies for CNNs, such as DBM pretraining with self-supervised learning (SSL) on DBM-processed T1 MRI data, model ensembling, domain-tailored attention, and few-shot learning techniques.
Ultimately, our results suggest that DBM uniquely captures subtle volumetric abnormalities across the brain and provides orthogonal information to the conventional features, contributing to MS progression in a way that is stable and clinically meaningful across multiple definitions of progression. Future work could integrate DBM with other MRI modalities used for lesion localization. Additionally, our study was restricted solely to patients with relapsing MS. Future work could utilize DBM to differentiate between different subtypes of MS or predict the conversion of relapsing MS to secondary progressive MS [75] when those subtypes are available. Our analysis pipeline is an initial starting point for more advanced data-driven studies that may provide better scientific and clinical insights about how MS impacts the brain.